# SymbolicAI: A framework for logic-based approaches combining generative models and solvers

### v2

@@ E.3 Translation
E.4 Filtering, Ranking, Extraction
@@
E.5 Implications
One of the main objectives behind developing SymbolicAI was to facilitate reasoning capabilities in conjunction with the statistical inference inherent in LLMs. Consequently, we can carry out deductive reasoning operations utilizing the Symbol objects. For instance, it is feasible to establish a series of operations with rules delineating the causal relationship between two symbols. The subsequent example illustrates the utilization of the & operator to compute the logical implication derived from the interaction of two distinct symbols.
@@ E.5 Implications
In the above example, the & operator overloads the and logical operator and extends its functionality. Furthermore, we can establish more sophisticated logical operators for and, or, and xor that can be grounded in formal proofs, as well as utilize the NeSy engine to parse data structures before evaluating the expressions. This enables the definition of bespoke operations for executing intricate and robust logical operations, incorporating constraints to validate outcomes and guide the computation towards the desired behavior.
E.6 Custom operations
@@
This is achieved using the _sym_return_type method, which can provide contextualized behavior based on the defined return type. More details can be found in the actual Symbol class.
In the context of LLMs, zero- and few-shot learning domains have emerged as essential techniques (Yao et al., 2023b; Shinn et al., 2023; Kim et al., 2023; Wei et al., 2022b; Lyu et al., 2023; Pitis et al., 2023; Madaan et al., 2023; Wang et al., 2022; Ye et al., 2023)666 This is by no means an exhaustive list, we only point the reader to some very influential and recent research. to enable models to generalize from limited training data and adapt to new tasks without requiring extensive retraining. This capability to learn and perform tasks with minimal examples is highly desirable, as it reduces the need for large labeled data sets and allows for faster deployment in new applications. In this section, we demonstrate how our Symbolic API incorporates Python decorators to define custom operations in the zero- and few-shot domains.
Consider the following example, where we define a custom operation to generate a random integer between 0 and 10 using the Symbolic API and Python decorators:
@@
The behavior of the prompt and constraints attributes is similar to the zero_shot decorator. The examples and limit arguments are new, with examples defining a list of instructions conditioning the NeSy engine, and limit specifying the maximum number of examples returned. The pre_processor and post_processor arguments accept lists of PreProcessor and PostProcessor objects, respectively, which are used to process the input before being fed into the NeSy engine and the output before being returned to the user. The wrp_kwargs argument passes additional arguments to the wrapped method, streamlining them towards the NeSy engine, or other engines.
E.7 Prompting
In this section, we discuss the design of prompts and their role in shaping the behavior of operations. Prompts serve as containers for information necessary to define specific operations, and the Prompt class serves as the base class for all the other Prompt classes. Consider the following example, where we define a Prompt for comparing two values using the NeSy engine. In it, when the <= operation on two Symbol objects will be resolved, the NeSy engine evaluates them in the context of the CompareValues prompt.
@@ E.7 Prompting
    Context-based Prompts are considered optional and can be defined in a static manner, either by sub-classing the Expression class and overwriting the static_context property, or at runtime by updating the dynamic_context property or passing a payload kwargs to a method. Below is an example of using the payload kwargs through the method signature:
@@
    Regardless of how the context is set, the contextualized prompt defines the desired behavior of Expression operations. For instance, if we want to operate in the context of a DSL without having to overwrite each base class method, we can use this approach777 See more details in this notebook..
@@
    Template Prompts are optional and encapsulate the resulting prediction to enforce a specific format. For example, to generate HTML tags, we can use a curated <html>...</html> template. This template enforces the NeSy engine to begin the generation process already in the context of an HTML tags format and not produce irrelevant descriptions about its task.
E.8 Complex expressions
@@
2  # very which case to use ‘self.value‘ contains the input
@@
The aforementioned example demonstrates the utilization of the causal_expression method, which allows us to extract information that can be resolved either manually or through external solvers, say WolframAlpha. Initially, when utilizing the GPT-3 backend, we anticipated a significant engineering effort to develop such a complex expression, as the GPT-3 backend frequently struggled with accurate information extraction and comparison resolution. However, we remained confident in the field’s progress, specifically with fine-tuned models like ChatGPT utilizing RLHF. We were delighted to witness these challenges being further tackled through the latest GPT-4 model (OpenAI, 2023).
@@
One of the most prominent illustrations of this concept is exhibited by Word2Vec (Mikolov et al., 2013). Word2Vec generates dense representations of words by training a shallow neural network to predict a word based on its neighboring words within a text corpus. These resulting vectors are extensively utilized in various natural language processing applications, including sentiment analysis, text classification, and clustering.
@@ 8  Compose(),
E.11 Streams
As demonstrated earlier, creating contextualized prompts refines the behavior of operations in the NeSy engine. However, this also consumes a considerable portion of the available context size. Given a limited context size, this constraint may pose challenges. Fortunately, the Stream processing expression offers a solution by opening a data stream and performing chunk-based operations on the input stream. Stream expressions can encapsulate other expressions. For instance, chunks can be managed using a Sequence expression, which permits multiple compositional operations in a sequential manner. The example below illustrates the definition of a Stream expression:
@@ 5    Outline(),
E.12 Error handling, debugging, and explainability
@@
E.13 Error handling
One of the fundamental aspects of the SymbolicAI API is being able to generate code. Consequently, errors may arise, and handling them contextually becomes vital. In pursuit of a self-evolving API, we introduce the Try expression, which includes built-in fallback statements and automatically retries execution after performing dedicated error analysis and correction. This expression analyzes both the input and the error, conditioning itself to resolve the error by modifying the original code888 This is similar to the recently released Auto-GPT application.. If the fallback expression succeeds, the result is returned; otherwise, the process is repeated for the specified number of retries. If the maximum number of retries is reached without resolving the issue, the error is raised again.
@@ E.13 Error handling
While not all errors can be resolved as easily as the demonstrated syntax error, we continue to explore more sophisticated error handling mechanisms, including the use of streams and clustering to address errors in a hierarchical and contextual manner.

### v3

@@ E.4 Translation
E.5 Filtering, Ranking, Extraction
@@
E.6 Implications
One of the main objectives behind developing SymbolicAI was to facilitate reasoning capabilities in conjunction with the statistical inference inherent in LLMs. Consequently, we can carry out deductive reasoning operations through the Symbol objects. For instance, it is feasible to establish a series of operations with rules delineating the causal relationship between two symbols. The subsequent example illustrates the utilization of the & operator to compute the logical implication derived from the interaction of two distinct symbols.
@@ E.6 Implications
In the above example, the & operator overloads the and logical operator and extends its functionality. Furthermore, we can establish more sophisticated logical operators for and, or, and xor that can be grounded in formal proofs and utilize the NeSy engine to parse data structures before evaluating the expressions. This enables the definition of bespoke operations for executing intricate and robust logical operations, incorporating constraints to validate outcomes and guide the computation towards the desired behavior.
E.7 Custom operations
@@
This is achieved through the _sym_return_type method, which can provide contextualized behavior based on the defined return type. More details can be found in the actual Symbol class.
In the context of LLMs, zero- and few-shot learning domains have emerged as essential techniques (Yao et al., 2023b; Shinn et al., 2023; Kim et al., 2023; Wei et al., 2022b; Lyu et al., 2023; Pitis et al., 2023; Madaan et al., 2023; Wang et al., 2022; Ye et al., 2023)777 This is by no means an exhaustive list, we only point the reader to some very influential and recent research. to enable models to generalize from limited training data and adapt to new tasks without requiring extensive retraining. This capability to learn and perform tasks with minimal examples is highly desirable, as it reduces the need for large labeled data sets and allows for faster deployment in new applications. In this section, we demonstrate how our Symbolic API incorporates Python decorators to define custom operations in the zero- and few-shot domains.
Consider the following example, where we define a custom operation to generate a random integer between 00 and $10$ using the Symbolic API and Python decorators:
@@
The behavior of the prompt and constraints attributes is similar to the zero_shot decorator. The examples and limit arguments are new, with examples defining a list of instructions conditioning the NeSy engine, and limit specifying the maximum number of examples returned. The pre_processor and post_processor arguments accept lists of PreProcessor and PostProcessor objects, respectively, which are utilized to process the input before being fed into the NeSy engine and the output before being returned to the user. The wrp_kwargs argument passes additional arguments to the wrapped method, streamlining them towards the NeSy engine, or other engines.
E.8 Prompting
In this section, we discuss the design of prompts and their role in shaping the behavior of operations. Prompts serve as containers for information necessary to define specific operations, and the Prompt class serves as the base class for all the other Prompt classes. Consider the following example, where we define a Prompt for comparing two values through the NeSy engine. In it, when the <= operation on two Symbol objects will be resolved, the NeSy engine evaluates them in the context of the CompareValues prompt.
@@ E.8 Prompting
    Context-based Prompts are considered optional and can be defined in a static manner, either by sub-classing the Expression class and overwriting the static_context property, or at runtime by updating the dynamic_context property or passing a payload kwargs to a method. Below is an example of payload kwargs through the method signature:
@@
    Regardless of how the context is set, the contextualized prompt defines the desired behavior of Expression operations. For instance, if we want to operate in the context of a DSL without having to overwrite each base class method, we can utilize this approach888 See more details in this notebook..
@@
    Template Prompts are optional and encapsulate the resulting prediction to enforce a specific format. For example, to generate HTML tags, we can utilize a curated <html>...</html> template. This template enforces the NeSy engine to begin the generation process already in the context of an HTML tags format and not produce irrelevant descriptions about its task.
E.9 Complex expressions
@@
The aforementioned example demonstrates the utilization of the causal_expression method, which allows us to extract information that can be resolved either manually or through external solvers, say WolframAlpha. Initially, when utilizing the GPT-3 backend, we anticipated a significant engineering effort to develop such a complex expression, as the GPT-3 backend frequently struggled with accurate information extraction and comparison resolution. However, we remained confident in the field’s progress, specifically with fine-tuned models like RLHF ChatGPT. We were delighted to witness these challenges being further tackled through the latest GPT-4 model (OpenAI, 2023).
@@
One of the most prominent illustrations of this concept is exhibited by Word2Vec (Mikolov et al., 2013a). Word2Vec generates dense representations of words by training a shallow neural network to predict a word based on its neighboring words within a text corpus. These resulting vectors are extensively utilized in various natural language processing applications, including sentiment analysis, text classification, and clustering.
@@ 8 Compose(),
E.12 Streams
As demonstrated earlier, creating contextualized prompts refines the behavior of operations in the NeSy engine. However, this also consumes a considerable portion of the available context size. Given a limited context size, this constraint may pose challenges. Fortunately, the Stream processing expression offers a solution by opening a data stream and performing chunk-based operations on the input stream. Stream expressions can encapsulate other expressions. For instance, chunks can be managed through a Sequence expression, which permits multiple compositional operations sequentially. The example below illustrates the definition of a Stream expression:
@@ 5 Outline(),
E.13 Error handling, debugging, and explainability
@@
E.14 Error handling
One of the fundamental aspects of the SymbolicAI API is being able to generate code. Consequently, errors may arise, and handling them contextually becomes vital. In pursuit of a self-evolving API, we introduce the Try expression, which includes built-in fallback statements and automatically retries execution after performing dedicated error analysis and correction. This expression analyzes both the input and the error, conditioning itself to resolve the error by modifying the original code999 This is similar to the recently released Auto-GPT application.. If the fallback expression succeeds, the result is returned; otherwise, the process is repeated for the specified number of retries. If the maximum number of retries is reached without resolving the issue, the error is raised again.
@@ E.14 Error handling
While not all errors can be resolved as easily as the demonstrated syntax error, we continue to explore more sophisticated error handling mechanisms, including streams and clustering to address errors in a hierarchical and contextual manner.
Appendix F Evaluation Details
In a computational graph, the VERTEX score compares the distribution of the generated model answer at each node against a reference distribution by sampling multiple valid trajectories at each node for the reference distribution and accounting for randomness through some predefined random trajectories. For instance, one of the predefined random trajectories in our benchmark was the string of ASCII characters which are considered printable, namely 0123456789abcdefghijklmnopqrstuvwxyzABCDEFGHIJKLMNOPQRSTUVWXYZ!"#\$%&’()*+, -./:;<=>?@[]_̂‘{|}~. Moreover, the VERTEX score is particularly well suited for the evaluation of multi-step workflows and in contexts where the solution space is or is expected to be diverse. We will now proceed by describing in detail the tasks that we defined in our benchmark.
F.1 Associative Prediction
We defined a total of 15 tasks involving in-context associations between two Symbol instances. SymbolicAI’s overloaded operators rely on predefined pseudo-grammars, as described in Section 4, that augment the operators with few-shot examples. For instance, the overloaded operator + utilized between two Symbol instances provides few-shot examples how to resolve additions with various data types:
1"’1’ + 2 =>3",
2"17 + pi =>20.1415926535...",
3"7.2 + five =>12.2",
4"True + 0 => False",
5"False + True =>False",
6"[’a’, b’] + [’c’, d’] =>[’a’, b’, c’, d’]",
7"[’apple’] + banana =>[’apple’, banana’]",
8"’Zero + 1 =>1",
9"’One + Two =>3",
10"’Three + 4 =>7",
11"’a + b + c + d =>a + b + c + d",
12...
Consequently, we can now test if the models can solve the addition between Symbol("two hundred and thirty four") and Symbol(7000).
F.2 Multi-modal Binding
We perform transformations between multiple modalities through language-based representations. Therefore, we need to evaluate the model’s proficiency in tool utilization, classification and routing of requests to relevant modules. We define a multi-modal Expression to detect the category of a task based on its content and to forward the task to the appropriate tool. The expression creates interfaces to tools like WolframAlpha for mathematical expressions, Selenium for website content scraping, SerpApi for search queries, and APILayer for optical character recognition. Each of the five tests aims to evaluate the appropriate handling of a specific type of input by the multi-modal Expression type, such as processing a website URL for scraping, interpreting a search engine query, testing if two vectors are linearly independent, comparing large numbers, and extracting text from an image. The following example shows the MultiModalExpression implementation of the forward function that uses isinstanceof operator on its own context to determine its current expression value and select the sub-routine that can evaluate the request.
1class MultiModalExpression(Expression):
2 def forward(self, ...):
3 formula = self.extract(’mathematical formula’)
4 ...
5 if self.isinstanceof(LINEAR_ALGEBRA):
6 ...
7 res = self.solver(formula)
8 res = res.query(’write a one sentence summary of the answer’)
9 ...
10 elif self.isinstanceof(NUMBER_COMPARISON):
11 res = self.solver(formula) # send directly to wolframalpha
12 else:
13 ...
14
15query = Symbol("is 100044347 bigger than 129981063.472?")
16expr = MultiModalExpression(query)
17res = expr(...)
F.3 Program Synthesis
We designed three separate tests related to program synthesis, where each task assesses the ability of the models to generate and execute code based on natural language instructions or provided templates:
1) The first task involves reading a LaTeX table template and data, then generating a function to populate the table with the given data.
2) The second task tests the automatic code generation for API calls by fetching data from a specified URL and extracting specific information from the retrieved content.
3) The third task evaluates the ability to construct a custom Expression that processes a Symbol through a specific Function component from the SymbolicAI package.
Each of the three tests follows a similar pattern, where the generated code is scored based on its similarity to valid references and normalized with random samples. Figure 13 shows possible samples from the third task category.
Figure 13: Samples for constructing custom expressions. The expected references are human-generated and are distributed according to the solution space. The set of valid references can be as well human-generated. The set of invalid (or undesirable) samples represent the subset we are the least interested in. Lastly, note that in general, the references need not be human-generated; any combination of human- and machine-generated references is possible. For instance, if synthetic data is utilized to distill knowledge from a larger model to a smaller model, the valid references and the expected references can be sampled from the larger model.
F.4 Logical Components
We designed six tests to assess the logical capabilities of the candidate models and group them as follows.
1) We utilize the Python library SymPy for symbolic mathematics to create the mathematical expression $ax+bx-cx-ay-by+cy+d$. The task for the model is then to factorize the expression and extract all unique symbols as a list.
1import sympy as sym
2
3...
4a, b, c, d, x, y = sym.symbols(’a, b, c, d, x, y’)
5expr = a * x + b * x - c * x - a * y - b * y + c * y + d
6# validate with sympy
7fact = sym.collect(expr, d, func=sym.factor)
8# model based factorization
9func = Factorization(’Factorize d from the expression such that your final start with: ‘d + (...‘:’)
10res = func(expr)
11# compare res with fact
12...
2) Three tasks evaluate a models’ capability to resolve the logical operations AND, OR, and XOR. For instance, the test for logical AND combines the symbols Symbol("The horn only sounds on Sundays") and Symbol("I hear the horn") and compares the answer against the human-generated references ”The horn only sounds on Sundays and I hear the horn.” and ”Since I hear the horn it is Sunday.” Since there is a large number of possible solutions, there is high variability in the solution space. Each model might prefer a different solution.
The following snippet shows how one can define a custom primitive class (CustomLogicPrimitive) for logic operators. The __or__ function gets overloaded and uses the built-in logic decorator from the core package to create a local function that evaluates two Symbol instances.
1from symai import core
2from symai.ops.primitives import Primitive
3
4class CustomLogicPrimitive(Primitive):
5 def __or__(self, other: Any) -> Any:
6 @core.logic(operator=’or’)
7 def _func(_, a: str, b: str):
8 pass # could impl. a fallback behavior here
9 return self._to_symbol(_func(self, other))
10 ...
11
12subject = ’cat’
13res = (Symbol(f’The {subject} has whiskers.’, primitives=CustomLogicPrimitive) | \
14 Symbol(f’The {subject} has a tail.’, primitives=CustomLogicPrimitive))
3) For another task we use a custom Expression that defines a DSL syntax and semantic structure. We use this Expression to extract higher-order logic expressions from a natural language statement, namely the puzzle ’Who is Jay’s brother?’101010 Bob has two sons, John and Jay. Jay has one brother and father. The father has two sons. Jay’s brother has a brother and a father. Who is Jay’s brother?, that preserves the original relationships. The following is a DSL snippet of the ’Who is Jay’s brother?’ puzzle:
1// Query
2IsBrotherOf(jay, john, bob) <- BrotherOf(jay, john) AND FatherOf(bob, jay) AND FatherOf(bob, john);
3
4// Facts
5BrotherOf(x, y) <- HAS(x, brother) AND HAS(y, brother) AND Sibling(x, y);
6FatherOf(x, y) <- HAS(x, son) AND ParentOf(x, y);
7ParentOf(x, y) <- IS(x, parent) AND IS(y, child);
8Sibling(x, y) <- IS(x, father) AND IS(y, father) OR IS(x, mother) AND IS(y, mother);
9
10...
4) For the final task, we again use the puzzle ’Who is Jay’s brother?’ to evaluate a models’ capability for complex conversions. We use the Z3 theorem prover (Moura & Bjørner, 2008) to solve the ’Who is Jay’s brother’ puzzle conditioned on the Z3 solvers’ solution to Einsteins’ famous puzzle ’Who owns the fish?’. The task involves an indirect translation from natural language to executable code by the Z3 solver; the solution to Einstein’s puzzle acts as a form of self-contained ”documentation” for how the Z3 solver should be utilized. The test constructs a template, which includes the task instructions, puzzle statement, and reference to the Einstein’s puzzle solution. The models are then asked to analyze the given problem and solution format and create a Python function with Z3 syntax that can solve the ’Who is Jay’s brother?’ puzzle. The dynamically generated code is executed within the test environment utilizing Python’s exec function. We check the access to the Z3 solver and run the generated solve_puzzle function supposed to contain the logic to solve the puzzle. Once executed, the assembled Z3 logical clauses are processed by the solver, which verifies that the set of constraints is satisfiable. If so, the model generated by the solver is queried for the puzzle’s solution and scored using our VERTEX score. The following is an example output from the Z3 representation of the solution to ’Who is Jay’s brother?’ puzzle:
1from z3 import Solver, Bool, And, Not, Const, BoolSort, EnumSort, Function, IntSort
2
3def solve_puzzle(S: Solver) -> Const:
4 # Define the enumeration sort for the individuals
5 Person, (BobE, JohnE, JayE, JaysBrotherE, FatherE) = EnumSort(’Person’, [’Bob’, ’John’, ’Jay’, ’JaysBrother’, ’Father’])
6
7 # Define a function from boolean to persons (for brother status)
8 is_brother = Function(’is_brother’, Person, BoolSort())
9
10 # Define the relationships
11 S.add(is_brother(JohnE) == True) # John is a brother
12 S.add(is_brother(JayE) == True) # Jay is a brother
13
14 ...
15
16 return query
F.5 Hierarchical Computational Graphs
In this section we extend on the hierarchical computational graphs section.
Research Paper Draft
The following example defines a Paper expression that takes in a sequence of expressions which are sequentially executed. The Method expression contains a Source expression which points to the actual human-written method. The Method expression acts as the root node that bootstraps the generation process. The RelatedWork expression contains a sequence of Cite expressions which are executed in parallel and utilized to define the context of the related work section. The Abstract and Title expressions get executed last because they require all the previous information to be available in their respective contexts. Each expression in the sequence of expressions from Paper takes in the context of its predecessors. All expressions also link their results to a global linker object which is utilized after the execution to retrieve individual results from the nodes of the expression’s computational graph. Each node was evaluated against its corresponding reference, all references representing actual sections from this research paper. The samples for each node were generated with a separate model (Claude 2) that was not part of this evaluation. In Figure 5 we show the resulting computational graph of the Paper expression.
1# define the computational graph
2expression = Paper(
3 Method(
4 # link to original code base where the main method is defined
5 Source(file_link=’/path/to/.../file’),
6 ),
7 # gather resources and write the related work
8 RelatedWork(
9 Cite(file_link=’/path/to/.../file’),
10 Cite(file_link=’/path/to/.../file’),
11 ...
12 ),
13 # write the abstract and title
14 Abstract(),
15 Title(),
16)
17# run the graph
18paper_result = expression(’Write a scientific paper’)
19# access linker to retreive the results from the method expression
20method = expr.linker.find(’Method’)
21# print result of the method expression
22print(method)
F.6 Caveats
One may have noticed that there are cases in which there is no need to sample multiple samples as there is only one expected answer, for instance, if we need to extract a specific number from a string and cast it to an integer. If the extraction process appends additional characters other than the number, the casting will fail. In such cases, the VERTEX score simply defaults to the chosen similarity measure, registering and penalizing any deviation from the expected answer.